OpenStudy (anonymous):

Solve the quadratic equation. 8x2 + 16x + 8 = 0

OpenStudy (anonymous):

8x^2 + 16x + 8 = 0 8 (x^2 + 2x + 1) = 0 8 (x+1) (x+1) = 0 Mathematically, 8 is not equal to 0. x+1 = 0 x = -1

OpenStudy (anonymous):

Did this help at all?